# Disentangling Representations through Multi-task Learning

### v2

@@ 5.1 Multi-task learning leads to disentangled representations
We train RNNs to do simultaneous classifications for $N_{\text{task}}$ linear partitions of the latent space for $D=2$ (Figure 2a, 6 partitions shown). To quantify the disentanglement of the representations after learning, we evaluate regression generalization by training a linear decoder to predict the ground truth $\mathbf{x}^{*}$ while network weights are frozen. We perform out-of-distribution 4-fold crossvalidation, i.e. train the decoder on 3 out of 4 quadrants and test in the remaining quadrant (Section A.2 for details). We also evaluate in-distribution (ID) performance by training the decoder in all quadrants. An example of train and test losses is shown in Figure S9f. We find that the network’s OOD and ID generalization performance are excellent (median $r^{2}=0.96,0.97$ respectively across 5 example networks); therefore the network has learned an abstract representation that zero-shot generalizes OOD. In addition, ID performance increases with the number of tasks $N_{\textit{task}}$, and the OOD generalization gap decreases (Figure 3a). Therefore we conclude that multi-task learning leads to abstract representations in the RNN’s hidden layer, when tasks span the latent space.
@@
Figure 3: Learning disentangled representations. (a) ID and OOD generalization performance for networks trained in different number of tasks $N_{\textit{task}}$. We report the 25, 50 and 75 percentile of $r^{2}$ for each network size (see Section A.2). ID and OOD performance increase with $N_{\textit{task}}$, and the generalization gap decreases, indicating that the networks have learned abstract representations. (b) The results hold for other autoregressive architectures, including LSTMs and GPT-2 transformers. (c) Angles between latent factor decoders (see Section A.3 for how they were estimated). The angles approach 90 degrees as $N_{\text{task}}\gg D$ for RNNs, but already fror $N_{\text{task}}\geq D$ for transformers. Remaining errors around 90 degrees are attributed to variability in the linear decoder fits. (d) Top 3 PCs of RNN activity ($N_{\textit{task}}=24$, $D=2$). Each line is a trial, while color saturation indicates time. All trials start from the center and move outwards, towards the location of $\mathbf{x}^{*}$ in state space. We color the last timepoint in each trial (squares) according to the quadrant this trial was drawn from. Red x’s correspond to attractors (see Section A.6). Here we remove input noise so that trajectories can be visualized easier. The network learns a two-dimensional continuous attractor that provides a disentangled representation of the 2D state space. (e) Spectral plot resulting from linearizing RNN dynamics around every fixed point (Section A.6). First two eigenvalues of the difference system are near 00, while the rest decay much faster, indicating marginal stability across two dimensions for every fixed point, a signature of a 2D continuous attractor.
@@
Since $\mathbf{x}^{*}$ can be decoded by this representation in unseen (by the decoder) parts of the state space, it follows that the representation can be used to solve any task involving the same latent variables, without requiring further pretraining. In other words, to solve any other task we do not need to deal with the denoising and unmixing of the latent factors $x_{1}$, $x_{2}$; we would just need to learn the (potentially non-linear) mapping from $x_{1}$, $x_{2}$ to task output. Furthermore, the representation scales linearly with input dimensionality $D$ (see Figure 5b). This marks a significant improvement from previously proposed models for representation learning in the brain where one task is executed at a time [49, 78], which scale linearly with $N_{\textit{task}}$, and exponentially with $D$ (see Section A.7 for details). Crucially, these findings are architecture-agnostic: they hold for non-leaky ("vanilla") RNNs, which outperform leaky ones for small $N_{\textit{task}}$, LSTMs which perform the best, and GPT-2 transformers (details in Section A.5) which have excellent performance already from $N_{\textit{task}}=2$ (Figure 3b). Note that state-space models have superior asymptotic performance, which is expected due to the nature of the task. We focus on leaky RNNs because of their closer correspondence to biological neurons, which have a membrane voltage that decays over time.
@@
So far we showcased abstractness, but not disentanglement. For disentaglement, it is crucial that the latents lie in orthogonal subspaces. Looking at the angles between the decoders of the latents, we find that they become orthogonal as $N_{\text{task}}\gg D$ for RNNs (Figure 3c), as predicted by our theory. Intriguingly, this already occurs from $N_{\text{task}}\geq D$ in transformers, showcasing their superior ability to separate latent factors. Furthermore, orthogonality strongly correlates with OOD generalization performance, which emphasizes the close link between disentanglement and abstractness: the more orthogonal the representations are, the cleaner the readout of the latent factors by linear decoders.
@@
We further demonstrate the robustness of our setting by showing that abstract representations emerge for different noise distributions and correlated inputs (Section A.8), non-linear boundaries (Section A.9), and for cognitive neuroscience integrate-to-bound tasks where the agent can make their decision whenever confident enough, not at a fixed time [40] (Section A.10).
@@ 5.2 Representational structure in RNNs and Transformers
In this section, we open the black box and investigate the representations learned by the networks, starting with RNNs. Figure 3d shows the top 3 PCs (capturing $\sim 85\,\%$ of the variance) of network activity after training (final accuracy $\sim 95\%$) for multiple trials, along with the fixed points of network dynamics. To find the fixed points, we follow a standard procedure outlined in Sussillo and Barak [70] (see Section A.6 for details). Looking at Figure 3d the fixed points span the entire two-dimensional manifold that the trials evolve in, which corresponds to a continuous attractor with stable states across a 2D "sheet". Linearizing the dynamics around each fixed point and computing the eigenvalues of the linearized system (Section A.6 for details), reveals marginal stability across two eigenvectors, i.e. near-00 eigenvalues which correspond to slow, integration dimensions in network dynamics, therefore confirming the continuousness of the attractor (Figure 3e). This implies that the network can store a short-term memory [75] of the current amount of accumulated evidence in a product space of the latent variables, and update it as further evidence arrives.
@@
Figure 4: RNN and GPT representations and relation to latent variables. (a) Hidden layer activations of RNN in Figure 2b (left) and GPT-2 transformer (right), while systematically varying the latent factors $x_{1}$ and $x_{2}$ from -0.5 to 0.5. Activations are plotted in 8*8 grids, one for each value of $x_{1}$ and $x_{2}$. Each grid contains firing rates for a total of 64 neurons for the RNN, and activations for 8 neurons for each of the 8 heads from the final embedding of the sequence for GPT-2. (b) Correlation coefficient of activations for both models with $x_{1}$ and $x_{2}$, respectively.
Finally, we examined RNN and GPT unit activations, and their relation to the latent variables. In Figure 4a we plot the steady-state firing rates for all 64 neurons for both networks, while regularly sampling $x_{1}$ and $x_{2}$. RNNs representations are sparse, with only $\sim 10\,\%$ of neurons active at any time, which is in line with sparse coding in the brain. In addition, the average firing rate is $\sim 1$ spike/s, which is surprisingly close to cortical values. Transformers on the other hand, do not have these features, shared by RNNs and their biological counterparts. Furthermore, we find that both networks display mixed selectivity, i.e. neurons are tuned to both variables, which is a known property of cortical neurons [63] (Figure 4b). This suggests that metrics of disentanglement that assume that individual neurons encode distinct factors of variation [30, 38, 13, 32, 21] might be insufficient in detecting disentanglement in networks that generalize well. While recent work incorporates such axis-alignment in the definition of disentaglement, our work along with others [37] showcases the advantages of approaching disentanglement from a mixed representations perspective. Importantly, these properties were not imposed during training, nor was there any parameter fine tuning involved; they emerged from task and optimization objectives.
@@ 5.3 Experiments confirm and extend theoretical predictions
An important prediction of our theory is that to learn abstract representations, $N_{\textit{task}}$ should exceed $D$. To test this, we increase $D$ (adding more inputs to Figure 2b), while varying $N_{\textit{task}}$. Sampling classification hyperplanes homogeneously (similar to Figure 2a, center) in high-dimensional spaces is non-trivial; therefore we resort to randomly sampling them. Figure 5b shows OOD generalization performance for various combinations of $D$ and $N_{\textit{task}}$. We observe that performance is bad when the $N_{\textit{task}}<D$, but it increases when $N_{\textit{task}}\geq D$. For RNNs, this increase is abrupt for smaller $D$ and more gradual for higher, which is in line with remarks by Johnston and Fusi [37] that it is easier to learn abstract representations when $D$ is high. Transformers on the other hand display higher generalization performance than RNNs, and always perform almost perfectly when $N_{\textit{task}}\geq D$, demonstrating their superior performance in learning abstract representations. Looking at the angles between latents for higher $D$ (Figure S10), we find that transformers have excellent disentanglement as long as $N_{\textit{task}}\geq D$, which might explain their superior generalization performance to RNNs for lower $N_{\textit{task}}$. These results, together with Figure 3c demonstrate the superior ability of transformers in disentangling latent factors. Overall, our findings confirm our theory that abstract representations emerge when $N_{\textit{task}}\geq D$, and even go beyond to suggest that disentangled representations emerge earlier than the theoretical condition $N_{\textit{task}}\gg D$, as long as the architecture is appropriate. These results are remarkable, especially for high $D$, because they go against our intuition that $N_{\textit{task}}$ should scale exponentially with $D$ to fill up the space adequately; instead it need only scale linearly.
@@ 6 Discussion
In this work, we proved that disentangled, generalizable representations must emerge in agents optimally solving multi-task evidence accumulation tasks canonical in the neuroscience literature. We also conducted experiments in a suite of autoregressive models (RNNs, LSTMs, transformers) which confirmed all of the main theoretical predictions. A key takeaway is that transformers more readily disentangle representations, which may explain their unique world understanding abilities.
@@ 6.1 Implications for representation learning
Our work has profound implications for learning representations that inherit the topological structure of the world. We prove this naturally happens as long as there are enough tasks to uniquely identify the location of $\mathbf{x}^{*}$. Crucially, the constraints from different tasks should be placed simultaneously on the representation, which explains why representations from context-dependent computation [49] are typically not disentangled.
@@
Our results provide a new perspective on the Platonic representation hypothesis [34], which suggests that the convergence in deep neural network representations is driven by a shared statistical model of reality, like Plato’s concept of an ideal “Platonic” reality. Theorem 3.1 suggests that the key factor driving convergence is the diversity and comprehensiveness of the tasks being learned. As long as individuals are faced with similar day-to-day tasks that collectively span the space of the underlying data representation, convergence to a shared, reality-aligned representation can occur. This could explain why for example modern LLMs come to encode high-level, human-interpretable concepts [71].
@@
While our problem is framed in terms of arbitrary injective observation map $f$, the formulation encompasses many scenarios relevant to the manifold hypothesis [23]. The function $f$ can represent a smooth manifold embedded in a high dimensional space, directly modelling the manifold hypothesis of deep learning. In neuroscience, $f$ could be a non-linear encoding of stimuli in a neural population response, connecting our work to neural manifold research [44]. By developing and testing theoretical guarantees for the emergence of disentangled representations in this multi-task problem formulation, we provide insight on how neural networks can inherently discover and linearize low-dimensional manifolds within high-dimensional, non-linear observations, enhancing our understanding of how complex data structures are captured and represented in deep learning models and biological systems alike.
@@
Finally, the theorem and experimental results provided here are not a one-way-street from dimensionality $D$ of the latents to how many tasks $N_{\text{task}}$ are required to uncover them. Instead, there is a fundamental interplay between richness of tasks performed and detail of the representation learned. In a high-dimensional world, the richness of the tasks at hand directly affects the dimensionality $D$ of the latents that can be extracted, allowing for "ground truths" $\mathbf{x}^{*}$ at different levels of granularity to be explored. The richer the label information available, the more fine-grained the resulting world model will be.
6.2 Limitations
A limitation of the present work is that factorization is assumed. Yet not all problems are factorizable, or should be factorized. For instance, a more coarse-grained understanding of the world, that doesn’t disentangle all factors, might be more suitable in many cases, and that might be reflected in the nature of the tasks. Furthermore, we focus on canonical cognitive neuroscience tasks which are somewhat removed from standard ML benchmarks. Normally, disentanglement methods would be tested against a benchmark such as dSprites [52]; however to the best of our knowledge no such benchmark exists for sequential tasks where evidence has to be aggregated over time. Future work could endeavor to apply our setting to richer tasks, like extracting latent item attributes from item embeddings when sequential decisions are made in online retailer settings.
@@ 6.3 Implications for neuroscience
The brain encodes variables of interest in a disentangled format, in processes as disparate as memory [8], emotion [67], and decision making [6]. Furthermore, performance in tasks has been shown to degrade once abstract representations collapse [67], supporting their role in guiding generalizable behavior. Given our findings, and that the cortical architecture is uniquely suited for parallel processing [28], the cortex is a prime candidate area for the construction of disentangled world models. Another such area is the thalamus; it is posited that thalamocortical loops operate in parallel, and combined with internal state-dependent mechanisms lead to state-dependent action selection (e.g. prioritizing water when thirsty), while evidence integration occurs in corticostriatal circuits [65]. The representations discovered here (continuous attractors, CANs) have been widely found in the brain when solving similar tasks, highlighting their role as a general computational substrate for cognitive functions in the brain (for relations of our work to the neuroscience literature, see Section A.11). The algorithmic efficiency of multi-task learning compared to alternatives (“context-dependent computation”, Mante et al. [49], Section A.7), makes us think that it is no coincidence that the cortex can support parallel processing; all the pieces are there, and we feel that the brain has to leverage this feature to construct faithful models of the world, as it does.
Acknowledgments and Disclosure of Funding
PV would like to thank the Onassis Foundation and AR the NOMIS Foundation for funding. AB thanks the NIH PTQN program for funding. No competing interests to declare. We would like to thank Yisong Yue for early discussions, Stefano Fusi for early feedback and Aiden Rosebush for discussions of proof methods.

### v3

@@ 5.1 Multi-task learning leads to disentangled representations
We train RNNs to do simultaneous classifications for $N_{\text{task}}$ linear partitions of the latent space for $D=2$ (Figure 2a, 6 partitions shown). To quantify the disentanglement of the representations after learning, we evaluate regression generalization by training a linear decoder to predict the ground truth $\mathbf{x}^{*}$ while network weights are frozen. We perform out-of-distribution 4-fold crossvalidation, i.e. train the decoder on 3 out of 4 quadrants and test in the remaining quadrant (Section A.2 for details). We also evaluate in-distribution (ID) performance by training the decoder in all quadrants. An example of train and test losses is shown in Figure S11f. We find that the network’s OOD and ID generalization performance are excellent (median $r^{2}=0.96,0.97$ respectively across 5 example networks); therefore the network has learned an abstract representation that zero-shot generalizes OOD. In addition, ID performance increases with the number of tasks $N_{\textit{task}}$, and the OOD generalization gap decreases (Figure 3a). Performance is identical when choosing a more nonlinear, power-law nonlinearity for the encoder (Section A.5). Therefore we conclude that multi-task learning leads to abstract representations in the RNN’s hidden layer, when tasks span the latent space.
@@
Figure 3: Learning disentangled representations. (a) ID and OOD generalization performance for networks trained in different number of tasks $N_{\textit{task}}$. We report the 25, 50 and 75 percentile of $r^{2}$ for each network size (see Section A.2). ID and OOD performance increase with $N_{\textit{task}}$, and the generalization gap decreases, indicating that the networks have learned abstract representations. (b) The results hold for other autoregressive architectures, including LSTMs and GPT-2 transformers. (c) Angles between latent factor decoders (see Section A.3 for how they were estimated). The angles approach 90 degrees as $N_{\text{task}}\gg D$ for RNNs, but already fror $N_{\text{task}}\geq D$ for transformers. Remaining errors around 90 degrees are attributed to variability in the linear decoder fits. (d) Top 3 PCs of RNN activity ($N_{\textit{task}}=24$, $D=2$), capturing 85% of variance (see inset). Each line is a trial, while color saturation indicates time. All trials start from the center and move outwards, towards the location of $\mathbf{x}^{*}$ in state space. We color the last timepoint in each trial (squares) according to the quadrant this trial was drawn from. Red x’s correspond to attractors (see Section A.7). Here we remove input noise so that trajectories can be visualized easier. The network learns a two-dimensional continuous attractor that provides a disentangled representation of the 2D state space. (e) Spectral plot resulting from linearizing RNN dynamics around every fixed point (Section A.7). First two eigenvalues of the difference system are near 00, while the rest decay much faster, indicating marginal stability across two dimensions for every fixed point, a signature of a 2D continuous attractor.
@@
Since $\mathbf{x}^{*}$ can be decoded by this representation in unseen (by the decoder) parts of the state space, it follows that the representation can be used to solve any task involving the same latent variables, without requiring further pretraining. In other words, to solve any other task we do not need to deal with the denoising and unmixing of the latent factors $x_{1}$, $x_{2}$; we would just need to learn the (potentially non-linear) mapping from $x_{1}$, $x_{2}$ to task output. Furthermore, the representation scales linearly with input dimensionality $D$ (see Figure 5b). This marks a significant improvement from previously proposed models for representation learning in the brain where one task is executed at a time (Mante et al., 2013; Yang et al., 2019), which scale linearly with $N_{\textit{task}}$, and exponentially with $D$ (see Section A.8 for details). Crucially, these findings are architecture-agnostic: they hold for non-leaky (”vanilla”) RNNs, which outperform leaky ones for small $N_{\textit{task}}$, LSTMs which perform the best, and GPT-2 transformers (details in Section A.6) which have excellent performance already from $N_{\textit{task}}=2$ (Figure 3b). Note that state-space models have superior asymptotic performance, which is expected due to the nature of the task. We focus on leaky RNNs because of their closer correspondence to biological neurons, which have a membrane voltage that decays over time.
@@
So far we showcased abstractness, but not disentanglement. For disentanglement, it is crucial that the latents lie in orthogonal subspaces. Looking at the angles between the decoders of the latents, we find that they become orthogonal as $N_{\text{task}}\gg D$ for RNNs (Figure 3c), as predicted by our theory. Intriguingly, this already occurs from $N_{\text{task}}\geq D$ in transformers, showcasing their superior ability to separate latent factors. Furthermore, orthogonality strongly correlates with OOD generalization performance, which emphasizes the close link between disentanglement and abstractness: the more orthogonal the representations are, the cleaner the readout of the latent factors by linear decoders.
@@
We further demonstrate the robustness of our setting by showing that abstract representations emerge for different noise distributions and correlated inputs (Section A.9), non-linear boundaries (Section A.10), and for cognitive neuroscience integrate-to-bound tasks where the agent can make their decision whenever confident enough, not at a fixed time (Krajbich et al., 2010) (Section A.11).
@@ 5.2 Representational structure in RNNs and Transformers
In this section, we open the black box and investigate the representations learned by the networks, starting with RNNs. Figure 3d shows the top 3 PCs (capturing $\sim 85\,\%$ of the variance) of network activity after training (final accuracy $\sim 95\%$) for multiple trials, along with the fixed points of network dynamics. To find the fixed points, we follow a standard procedure outlined in Sussillo & Barak (2013) (see Section A.7 for details). Looking at Figure 3d the fixed points span the entire two-dimensional manifold that the trials evolve in, which corresponds to a continuous attractor with stable states across a 2D ”sheet”. Linearizing the dynamics around each fixed point and computing the eigenvalues of the linearized system (Section A.7 for details), reveals marginal stability across two eigenvectors, i.e. near-00 eigenvalues which correspond to slow, integration dimensions in network dynamics, therefore confirming the continuousness of the attractor (Figure 3e). This implies that the network can store a short-term memory (Wang, 2001) of the current amount of accumulated evidence in a product space of the latent variables, and update it as further evidence arrives.
@@
Figure 4: RNN and GPT representations and relation to latent variables. (a) Hidden layer activations of RNN in Figure 2b (left) and GPT-2 transformer (right), while systematically varying the latent factors $x_{1}$ and $x_{2}$ from -0.5 to 0.5. Activations are plotted in 8*8 grids, one for each value of $x_{1}$ and $x_{2}$. Each grid contains firing rates for a total of 64 neurons for the RNN, and activations for 8 units for each of the 8 heads from the final embedding of the sequence for GPT-2. (b) Correlation coefficient of activations for both models with $x_{1}$ and $x_{2}$, respectively.
Finally, we examined RNN and GPT unit activations, and their relation to the latent variables. In Figure 4a we plot activations for all 64 units for both networks, while regularly sampling $x_{1}$ and $x_{2}$. RNNs representations are sparse, with only $\sim 10\,\%$ of neurons active at any time, which is in line with sparse coding in the brain (see Section A.12 for quantification of sparsity as a function of $N_{\textit{task}}$, $D$ and RNN architecture). In addition, the average firing rate is $\sim 1$ spike/s, which is surprisingly close to cortical values. Transformers on the other hand, do not have these features, shared by RNNs and their biological counterparts. Furthermore, we find that both networks display mixed selectivity, i.e. neurons are tuned to both variables, which is a known property of cortical neurons (Rigotti et al., 2013) (Figure 4b). This suggests that metrics of disentanglement that assume that individual neurons encode distinct factors of variation (Higgins et al., 2017; Kim & Mnih, 2018; Chen et al., 2018; Eastwood et al., 2022; Hsu et al., 2023) might be insufficient in detecting disentanglement in networks that generalize well. While recent work incorporates such axis-alignment in the definition of disentaglement, our work along with others (Johnston & Fusi, 2023) showcases the advantages of approaching disentanglement from a mixed representations perspective. Importantly, these properties were not imposed during training, nor was there any parameter fine tuning involved; they emerged from task and optimization objectives.
@@ 5.3 Experiments confirm and extend theoretical predictions
An important prediction of our theory is that to learn abstract representations, $N_{\textit{task}}$ should exceed $D$. To test this, we increase $D$ (adding more inputs to Figure 2b), while varying $N_{\textit{task}}$. Sampling classification hyperplanes homogeneously (similar to Figure 2a, center) in high-dimensional spaces is non-trivial; therefore we resort to randomly sampling them. Figure 5b shows OOD generalization performance for various combinations of $D$ and $N_{\textit{task}}$. We observe that performance is bad when the $N_{\textit{task}}<D$, but it increases when $N_{\textit{task}}\geq D$. For RNNs, this increase is abrupt for smaller $D$ and more gradual for higher, which is in line with remarks by Johnston & Fusi (2023) that it is easier to learn abstract representations when $D$ is high. Transformers on the other hand display higher generalization performance than RNNs, and always perform almost perfectly when $N_{\textit{task}}\geq D$, demonstrating their superior performance in learning abstract representations. Looking at the angles between latents for higher $D$ (Figure S12), we find that transformers have excellent disentanglement as long as $N_{\textit{task}}\geq D$, which might explain their superior generalization performance to RNNs for lower $N_{\textit{task}}$. These results, together with Figure 3c demonstrate the superior ability of transformers in disentangling latent factors. Overall, our findings confirm our theory that abstract representations emerge when $N_{\textit{task}}\geq D$, and even go beyond to suggest that disentangled representations emerge earlier than the theoretical condition $N_{\textit{task}}\gg D$, as long as the architecture is appropriate. These results are remarkable, especially for high $D$, because they go against our intuition that $N_{\textit{task}}$ should scale exponentially with $D$ to fill up the space adequately; instead it need only scale linearly.
@@ 6 Discussion
In this work, we proved that disentangled, generalizable representations must emerge in agents optimally solving multi-task evidence accumulation tasks canonical in the neuroscience literature. We also conducted experiments in a suite of autoregressive models (RNNs, LSTMs, transformers) which confirmed all of the main theoretical predictions. A key takeaway is that transformers more readily disentangle representations, which may explain their unique world understanding abilities. Here we discuss the broader impact of this work for representation learning and neuroscience alike. Limitations of this study and how it can be extended in the future are discussed in Section A.15.
@@ 6.1 Implications for representation learning
Our work has profound implications for learning representations that inherit the topological structure of the world. We prove this naturally happens as long as there are enough tasks to uniquely identify the location of $\mathbf{x}^{*}$. Crucially, the constraints from different tasks should be placed simultaneously on the representation, which explains why representations from context-dependent computation (Mante et al., 2013) are typically not disentangled.
@@
Our results provide a new perspective on the Platonic representation hypothesis (Huh et al., 2024), which suggests that the convergence in deep neural network representations is driven by a shared statistical model of reality, like Plato’s concept of an ideal “Platonic” reality. Theorem 3.1 suggests that the key factor driving convergence is the diversity and comprehensiveness of the tasks being learned. As long as individuals are faced with similar day-to-day tasks that collectively span the space of the underlying data representation, convergence to a shared, reality-aligned representation can occur. This could explain why for example modern LLMs come to encode high-level, human-interpretable concepts (Templeton et al., 2024).
@@
While our problem is framed in terms of arbitrary injective observation map $f$, the formulation encompasses many scenarios relevant to the manifold hypothesis (Fefferman et al., 2013). The function $f$ can represent a smooth manifold embedded in a high dimensional space, directly modelling the manifold hypothesis of deep learning. In neuroscience, $f$ could be a non-linear encoding of stimuli in a neural population response, connecting our work to neural manifold research (Langdon et al., 2023). By developing and testing theoretical guarantees for the emergence of disentangled representations in this multi-task problem formulation, we provide insight on how neural networks can inherently discover and linearize low-dimensional manifolds within high-dimensional, non-linear observations, enhancing our understanding of how complex data structures are captured and represented in deep learning models and biological systems alike.
@@
Finally, the theorem and experimental results presented here are not a one-way-street from dimensionality $D$ of the latents to how many tasks $N_{\text{task}}$ are required to uncover them. Instead, there is a fundamental interplay between richness of tasks performed and detail of the representation learned. In a high-dimensional world, the richness of the tasks at hand directly affects the dimensionality $D$ of the latents that can be extracted, allowing for ”ground truths” $\mathbf{x}^{*}$ at different levels of granularity to be explored. The richer the label information available, the more fine-grained the resulting world model will be.
@@ 6.2 Implications for neuroscience
The brain encodes variables of interest in a disentangled format, in processes as disparate as memory (Boyle et al., 2022), emotion (Saez et al., 2015), and decision making (Bongioanni et al., 2021). Furthermore, performance in tasks has been shown to degrade once abstract representations collapse (Saez et al., 2015), supporting their role in guiding generalizable behavior. Our findings put forth parallel processing as a unifying mechanism for generalization in brains. The cortex, with its massively parallel architecture (Markram et al., 2015; Hawkins et al., 2019), is a prime candidate area for the construction of disentangled, generalizable world models. Another candidate area is the thalamus; it is posited that thalamocortical loops operate in parallel, and combined with internal state-dependent mechanisms lead to state-dependent action selection (e.g. prioritizing water when thirsty), while evidence integration occurs in corticostriatal circuits (Rubin et al., 2020). The representations discovered here (continuous attractors, CANs) have been widely found in the brain when solving similar tasks, highlighting their role as a general computational substrate for cognitive functions in the brain (for relations of our work to the neuroscience literature, see Section A.13). Notably, the receipt of rich supervisory signals from the environment is not a requirement for our setting, as it can leverage the output of previously learned tasks (see Section A.14 on the biological plausibility of multi-task learning).
The algorithmic efficiency of multi-task learning compared to alternatives (“context-dependent computation”( Mante et al. (2013), Section A.8)), makes us think that it is no coincidence that the cortex can support parallel processing; all the pieces are there, and we feel that the brain has to leverage this feature to construct faithful models of the world, as it does.
Reproducibility Statement
All code used to generate the results can be found in https://github.com/panvaf/DisentangleRes. The experiments are seeded, ensuring exact reproducibility of results. Five networks have been trained for every configuration shown, to provide sufficient statistics to support our conclusions. To ensure full clarity of the theoretical results in the main text, a full proof is provided in Appendix B.
Author Contributions
P.V. conceived the project, wrote the code, performed the experiments and wrote the manuscript. A.B. developed the theory, wrote the theory sections and parts of the manuscript. A.R. supervised the research and provided funding.
Acknowledgments
P.V. would like to thank the Onassis Foundation and A.R. the NOMIS Foundation for funding. A.B. thanks the NIH PTQN program for funding. No competing interests to declare. We would like to thank Yisong Yue for early discussions, Aiden Rosebush for discussions of proof methods and Stefano Fusi for early feedback.